# ImageNot: A contrast with ImageNet preserves model rankings

### v1

@@ Abstract
We introduce ImageNot, a dataset designed to match the scale of ImageNet while differing drastically in other aspects. We show that key model architectures developed for ImageNet over the years rank identically when trained and evaluated on ImageNot to how they rank on ImageNet. This is true when training models from scratch or fine-tuning them. Moreover, the relative improvements of each model over earlier models strongly correlate in both datasets. We further give evidence that ImageNot has a similar utility as ImageNet for transfer learning purposes. Our work demonstrates a surprising degree of external validity in the relative performance of image classification models. This stands in contrast with absolute accuracy numbers that typically drop sharply even under small changes to a dataset.
@@ 1 Introduction
The Achilles heel of machine learning is its apparent lack of robustness. Models developed in one domain consistently perform worse in new domains. A slew of important—but essentially negative—empirical results demonstrated that even small changes to a dataset can impact the accuracy of machine learning models dramatically (Recht et al., 2019; Taori et al., 2020; Miller et al., 2021). There is good reason to fear that machine learning lacks external validity (Torralba and Efros, 2011; Liao et al., 2021).
@@
Against this pessimistic backdrop, we show that key qualities of machine learning models do generalize to a surprising degree. The starting point of our work is a thought experiment. How different a dataset from ImageNet would have supported essentially the same model developments over the years? ImageNet was at the center of the deep learning revolution in computer vision of the 2010s. Numerous deep architectures were developed specifically with ImageNet in mind. It’s, therefore, reasonable to conjecture that a different benchmark would have spurred different model developments. Our investigation suggests otherwise.
@@
To argue this point, we create a contrast with ImageNet, a new dataset called ImageNot. ImageNot has 1000 classes with 1000 examples per class, just like ImageNet’s ILSVRC-2012 distribution, but ImageNot differs in key characteristics associated with ILSCRC-2012 (Russakovsky et al., 2015; Deng et al., 2009). No human annotators provided any labels. ImageNot pictures instead come from noisy web-crawled data selected based on the image captions and class names alone. While the classes in ILSVRC-2012 have strong concept overlap, the ones we chose for ImageNot are deliberately arbitrary and unrelated. Given the concerns about model architectures overfitting to ImageNet, one expects a new dataset like ImageNot designed to contrast ImageNet to give a range of results that diverge from those associated with ImageNet. Surprisingly, ImageNot turns out to share many of ImageNet’s qualities in terms of model training, comparison and ranking, fine-tuning, and transfer learning.
@@ 1.1 Our contributions
(a) Model rankings with models initialized with random weights, trained on a dataset and evaluated on a held-out test set. We note that on CIFAR10, the error bars for EfficientNet-V2-L (Tan and Le, 2019) and ConvNext-L (Liu et al., 2022) overlap, so the rankings may still be exactly preserved.
(b) Relative progress of each model architecture. For a set of test accuracies $X$, relative progress is $X_{i}/X_{\text{AlexNet}}$. This figure is reproduced in Figure 8(b).
Figure 1: Final Model Ranking and Relative Improvement.
@@
Our primary contribution is an in-depth investigation into the external validity, robustness, and reproducibility of model development progress in machine learning. Our work spans the creation of a new dataset and the compute-intensive training and evaluation of numerous architectures on the new dataset.
Creating ImageNot.
@@
We create the ImageNot dataset from the LAION-5B dataset, a public collection of 5 billion image-text pairs crawled from the web and filtered by the CLIP model (Schuhmann et al., 2022; Radford et al., 2021). Although CLIP is a model trained with knowledge of ImageNet, the sheer size of LAION-5B ensures that images in LAION-5B are not already too similar to ImageNet images. We choose 1000 classes from the WordNet lexical database while avoiding nouns close to those included in ImageNet. For each class, we pick 1000 images from LAION-5b based on caption-class similarity alone. We use Roberta text embeddings for this purpose. As a text-only model, Roberta was never trained on ImageNet. This mitigates the concern that our selection of images from LAION-5B might be biased in favor of ImageNet-like instances. We give a detailed analysis of the differences between ImageNet and ImageNot.
ImageNot preserves model rankings and relative improvements.
@@
Our primary empirical finding is that ImageNot has the same image classification model rankings as ImageNet. This finding is important because it provides evidence that the model architecture improvements centered around ImageNet could have occurred similarly for a different dataset than ImageNet. We consider 6 representative model architectures from AlexNet to ConvNeXt-L (Krizhevsky et al., 2012; Liu et al., 2022). We train from scratch and evaluate each model on ImageNot. We observe that the final ranking is the same as on ImageNet. Moreover, the improvement that each model makes relative to prior models is roughly the same on ImageNot as it is on ImageNet. Figure 1 summarizes these results. The remarkable consistency of relative improvements suggests an intriguing counterfactual. Had ImageNot been the benchmark at the time of a new image classification model development, it could have induced the same improvements as ImageNet.
@@
ImageNot pretraining and transfer learning.
ImageNet has proved useful for pretraining and transfer learning (Kornblith et al., 2019). We wanted to know if the same is true for ImageNot. Toward this end, we conducted two experiments for each model:
@@
    Pretrain on ImageN[o/e]t, finetune all layers on CIFAR-10, evaluate on CIFAR-10.
@@
    Pretrain on ImageN[o/e]t, finetune last layer on CIFAR-10, evaluate on CIFAR-10.
We find that a model’s test accuracy on ImageN[o/e]t is a good predictor of fine-tunability and transferability. These experiments suggest that ImageNot shares with ImageNet its utility for pretraining and transfer learning.
To summarize, rankings of image classification models are surprisingly stable across significantly different datasets. Prior work had observed that small variations in a data-generating process preserve model rankings. Our work adds a more extreme contrast. Even significant deliberate deviations from a dataset can maintain essential characteristics of model training and comparison. In particular, our findings suggest that the external validity of machine learning is, in important ways, significantly greater than previously assumed.
Code.
@@ 2 Discussion of related work
The external validity of benchmark results for image classification has long been of concern. Indeed, benchmarks often contain biases that do not generalize to the data observed in the real world. Ponce et al. (2006) provide evidence of the narrow range of variation within existing benchmarks in the early stages of image classification research. Subsequently, Torralba and Efros (2011) compare multiple well-known image classification datasets, including SUN09 (Xiao et al., 2010), PASCAL Visual Object Classes (VOC) (Everingham et al., 2010), and Caltech-101 (Fei-Fei et al., 2004). They find evidence of selection bias, capture bias, and negative set bias in these datasets.
The desire for a more representative benchmark for real-world image classification tasks gave birth to ImageNet, a large-scale image database (Deng et al., 2009). The extensive scale and relatively meticulous curation of ImageNet were thought to alleviate some of the limitations identified in earlier datasets, launching over a decade of active development in image classification, largely centered around improving accuracy on the ImageNet dataset. Still, there are concerns about the external validity of ImageNet results. Tsipras et al. (2020) studied the design choices of ImageNet and their effect on alignment between the ImageNet dataset and the broader underlying object recognition task. They concluded that ImageNet contains some inherent biases that cause misalignment. Despite this misalignment, they argue, improvements in multiclass classification on ImageNet still translate to the underlying object recognition task. However, they raised the issue that these biases make it difficult to identify actual improvements in object recognition. Xiao et al. (2020) identify one of these biases in that models tend to rely on the background of ImageNet images, which are correlated with the corresponding labels.
Recht et al. (2019) investigated whether ImageNet classifiers generalize to a newly curated ImageNet test set—recall that by ImageNet, we generally refer to the ILSVRC 2012 ImageNet challenge. They found that accuracies decrease strongly from the original the new test set, in the range of 11%-14%. Despite these findings, Recht et al. (2019)’s results showed that model rankings did not significantly change between the two ImageNet test sets, meaning, if model a performed better than model b on the original ImageNet test set, it also performed better on the new ImageNet test set. This speaks to the stability of model rankings under a slight distribution shift (new versus old test set) but leaves the question of how stable model rankings are under more drastic changes to the dataset. Kornblith et al. (2019) showed that for a set of popular image classification architectures, improvements on ImageNet are often a good predictor of transfer learning performance on other image classification datasets. In addition, beyond image classification, Huang et al. (2017b) showed that improvements in image classification on ImageNet often also predicted improvements in other computer vision tasks, such as image segmentation.
While considerable research on the generalizability of ImageNet models has been centered around changes in accuracy under distribution shift, this work proposes a different perspective. We argue that for benchmarking purposes, providing robust model rankings is a more critical property of an effective benchmark than preserving accuracy or similar metrics. This, however, does not discount the importance of metrics such as accuracy on a well-thought-out and well-curated dataset.
In Mark Liberman’s seminal 2015 lecture on reproducible research and the common task method, he described the role of benchmarks as an aligning mechanism for measuring progress (Liberman, 2015). We distinguish between two related functions of benchmarks in measuring progress: (i) Ranking, i.e., measuring relative progress, (ii) quantifying, i.e., absolute metrics such as accuracy in image classification. Although metrics can be useful, we argue they can be problematic targets for gauging external validity. Consider two identical benchmarks where we inserted a different fraction of random labels in each benchmark. The accuracy numbers may, therefore, be completely different, while the two benchmarks fundamentally measure the same thing. We, therefore, focus on ranking as the primary gauge of external validity.
@@
The observed progress on image classification can be well summarized by a set of models that span the active development of model architectures around ImageNet: AlexNet (Krizhevsky et al., 2012), VGG (Simonyan and Zisserman, 2014), ResNet (He et al., 2016), DenseNet (Huang et al., 2017a), EfficientNet (Tan and Le, 2019), and ConvNeXt (Liu et al., 2022). While other models were also developed during the so-called ImageNet era, we focus on models that were most impactful and represented a notable shift in architecture design – we include a discussion in Appendix A. Specifically, we considered the availability of model implementations across hubs such as TorchHub (PyTorch Team, 2024), TensorFlow (Abadi et al., 2015), and Huggingface (Wolf et al., 2020) to be an indicator of popularity and, therefore, impact. We note that while newer architectures such as Vision Transformers  (Devlin et al., 2019; Dosovitskiy et al., 2020) are now commonplace in the literature, we only consider architectures developed to improve ImageNet accuracy in our study.
Many variants of ImageNet have been created to test the robustness of ImageNet models. Hendrycks and Dietterich (2019) introduced ImageNet-C/P, ImageNet with common corruptions.  Hendrycks et al. (2020) introduced ImageNet-R, stylized distribution shifts on ImageNet. Li et al. (2022) introduced ImageNet-W, ImageNet with watermarks, a learning shortcut. Shirali and Hardt (2023) introduced LAIONet, an intended recreation of ImageNet from the LAION-400M dataset, in order to identify key differences between ImageNet and LAIONet.
Our aim is distinctly different from these works. Rather than studying variants of ImageNet, we create an entirely different dataset. We aim to evaluate if this completely different dataset suggests the same model development trajectory, i.e., incremental rankings, as ImageNet. Accordingly, we introduce a new dataset, ImageNot, with entirely different classes and a unique curation process, drawing inspiration from LAIONet’s data-generating process (Shirali and Hardt, 2023). A distinct difference between our work and these previous works is that we train and test on this new dataset. This contrasts to (Recht et al., 2019; Shirali and Hardt, 2023), who compare test accuracies of the same ImageNet trained models on two different test sets with ImageNet classes. Our results compare the test accuracy of ImageNet-trained models on ImageNet’s test set with the test accuracy of ImageNot-trained models on ImageNot. This removes the confound of ImageNet training and the ImageNet classes, better isolating the external validity of model architectures independent of their training data. Additionally, we developed ImageNot rather than other datasets to ensure the strength of its contrast with ImageNet.
@@ 3 Constructing ImageNot
Figure 2: ImageNot caption x synset-gloss RoBERTa embeddings similarity for each class. We show the minimum, mean, median, and maximum similarity distribution across the 1000 classes.
@@
We develop ImageNot, a dataset of the same scale as ImageNet’s ILSVRC-2012 – 1000 classes with 1000 examples per class – but without key design choices that have been thought critical in ImageNet’s effectiveness (Russakovsky et al., 2015). First, ImageNot has deliberately arbitrary classes constrained only to be strictly different from those in ImageNet. Second, we do not utilize any human annotators. There were no humans in the loop in selecting specific image and label pairs. We create the ImageNot dataset from the LAION-5B dataset, an academic, public massive scale collection of 5 billion image-text pairs crawled from the web and filtered by the CLIP model (Schuhmann et al., 2022; Radford et al., 2021). This is in contrast with the ILSVRC 2012 challenge, a subset of the ImageNet database which was constructed from Flickr image search results (Deng et al., 2009).
Similar to ImageNet, ImageNot classes correspond to nouns in WordNet synsets found in LAION-5B captions, where WordNet is a large lexical database of English (Fellbaum, 1998; Miller, 1995; Hardeniya et al., 2016). The following definitions are important for describing WordNet. A given word has a set of synsets, where a synset is a sense of a word and has a corresponding short definition called a gloss. For example, ‘car’ has 2 synsets with corresponding glosses: (i) “a motor vehicle with four wheels; usually propelled by an internal combustion engine” and (ii) “a wheeled vehicle adapted to the rails of a railroad.” Additionally, each synset consists of a set of synonyms called lemmas. The lemmas of ‘car’-(i) include automobile, motorcar, while the lemmas of ‘car’-(ii) include railcar, railwaycar. A WordNet tree represents a semantic hierarchy of synsets via an ‘IS-A’ structure, i.e., a synset is a ‘type’ of its synset parent, e.g., mammal $\rightarrow$ placental $\rightarrow$ carnivore $\rightarrow$ dog $\rightarrow$ working dog $\rightarrow$ husky.
Each class in ImageNot represents a unique noun synset. At a high level, for each class, we select images from LAION-5B whose captions are similar to the synset corresponding to the class. We utilize RoBERTa (Liu et al., 2019) embeddings, an optimized version of BERT (Devlin et al., 2019), to estimate similarity between text sequences. Since RoBERTa was never trained on ImageNet, this selection mechanism mitigates the concern that our image selection process may be biased in favor of ImageNet-like images.
We note that creators of LAION-5B’s only included image and caption pairs that are sufficiently similar based on CLIP embeddings (Contrastive Language–Image Pre-training (Radford et al., 2021)), a multimodal deep learning model (Schuhmann et al., 2022).
Constructing ImageNot.
To choose the 1000 synsets (classes) that are included in ImageNot, we do the following (employing similar mechanics as Shirali and Hardt (2023)):
@@
    Filter out synsets considered ‘Not Safe for Work’ (NSFW) by Song et al. (2021)’s NSFW classifier.
@@
    Filter out synsets with an average top n caption x synset-gloss pair RoBERTa embeddings similarity (normalized inner product) less than 0.82  (Liu et al., 2019). We found 0.82 to be the highest reasonable choice to allow for sufficient examples in enough classes (Figure 3).
@@
    Finally, select the top 1000 classes ordered by the average similarity between caption x synset-gloss pairs (Liu et al., 2019) (desideratum 4).
Finally, we obtain ImageNot, a large-scale dataset that contrasts ImageNet. ImageNot has an average caption-label definition similarity of 0.91. Next, we compare ImageNot to ImageNet. ImageNot classes can be found at https://github.com/olawalesalaudeen/imagenot.
@@
In this section, we delve into the differences between ImageNot and ImageNet, focusing on their design aspects. The primary distinctions lie in two areas: (i) the classes in ImageNot are arbitrarily different from those in ImageNet, and (ii) the attribution of a synset to an image in ImageNot is conducted without human intervention. We begin by examining the key differences between the two datasets.
The ‘Not’ in ImageNot.
@@
The first significant not is the hierarchy of synsets, based on the WordNet tree. ImageNet is constructed with synsets representing a densely populated subset of the WordNet tree. For instance, ImageNet contains 147 categories of dogs. In contrast, ImageNot is not designed to have densely populated subsets of the WordNet tree. Moreover, ImageNot synsets tend to be higher up in the WordNet tree (Figure 4) with ImageNot synsets having a median depth of 8 while ImageNet synsets have a median depth of 10 (more details about the WordNet subtrees can be found in Appendix B.1 Table 3). One consequence of this difference in median depth is that the average diversity of images in an ImageNot synset is higher than that of ImageNet. We illustrate the distribution of intra-class similarity within each dataset in Figure 6, where ImageNet exhibits higher intra-class diversity than ImageNot.

### v2

@@ Abstract
We introduce ImageNot, a dataset constructed explicitly to be drastically different than ImageNet while matching its scale. ImageNot is designed to test the external validity of deep learning progress on ImageNet. We show that key model architectures developed for ImageNet over the years rank identically to how they rank on ImageNet when trained from scratch and evaluated on ImageNot. Moreover, the relative improvements of each model over earlier models strongly correlate in both datasets. Our work demonstrates a surprising degree of external validity in the relative performance of image classification models when trained and evaluated on an entirely different dataset. This stands in contrast with absolute accuracy numbers that typically drop sharply even under small changes to a dataset.
@@ 1 Introduction
The Achilles heel of machine learning is its apparent lack of robustness. Models developed in one domain consistently perform worse in new domains. A slew of important—but essentially negative—empirical results demonstrated that even small changes to a dataset can impact the accuracy of machine learning models dramatically (recht2019imagenet; taori2020measuring; miller2021accuracy). Consequently, there is good reason to fear that machine learning lacks external validity (torralba2011unbiased; liao2021we).
@@
Against this pessimistic backdrop, we show that key qualities of machine learning models do generalize to a surprising degree. The starting point of our work is a thought experiment. How different a dataset from ImageNet would have supported essentially the same model developments over the years? ImageNet was at the center of the deep learning revolution in computer vision of the 2010s. Numerous deep architectures were developed specifically with ImageNet in mind. It’s therefore reasonable to conjecture that a different benchmark would have spurred different model developments. Our investigation suggests otherwise.
@@
To argue this point, we create a contrast with ImageNet, a new dataset called ImageNot. ImageNot has 1000 classes with 1000 examples per class, just like ImageNet’s ILSVRC-2012 distribution, but ImageNot differs in key characteristics associated with ILSVRC-2012 (ILSVRC15; deng2009imagenet). No human annotators provided any labels. ImageNot pictures instead come from noisy web-crawled data selected based on the image captions and class names alone. While the classes in ILSVRC-2012 have strong concept overlap, the ones we chose for ImageNot are deliberately arbitrary and unrelated.
The central question we ask is: Does ImageNot provide the same model rankings as ImageNet? An important difference between our work and previous work is that they investigate the robustness of ImageNet models, models trained on ImageNet. In contrast, we investigate the external validity of the core deep learning developments on ImageNet; thus, we remove the ImageNet confound, unlike in previous work. The architectures developed on ImageNet are trained from scratch and evaluated on ImageNot. More generally, we identify five major confounds that hinder interpretations of previous work in the same context of this work on external validity: Previous work had (i) evaluation sets with same images as ImageNet but with corruptions (Hendrycks2019BenchmarkingNN; hendrycks2021natural; Hendrycks2020TheMF), (ii) the same data sources, Flickr (recht2019imagenet), (iii) the same image classes (recht2019imagenet; barbu2019objectnet; Hendrycks2019BenchmarkingNN; hendrycks2021natural; Hendrycks2020TheMF), and (iv) the same data generative mechanisms, i.e., the procedure of human annotation (recht2019imagenet). Furthermore, (v) the same models trained on ImageNet are evaluated (recht2019imagenet; barbu2019objectnet; Hendrycks2019BenchmarkingNN; hendrycks2021natural; Hendrycks2020TheMF). While the results of these previous works have been interpreted in the context of external validity of deep learning model development, i.e., scientific progress on ImageNet, these confounds limit the soundness of such interpretations, since they may be responsible for the observed robustness of rankings. These confounds leave the question of the external validity of ImageNet as a benchmark for model development progress unanswered. Importantly, the goal of this work is not to propose ImageNot as a new general benchmark or replacement for ImageNet. ImageNot is a construct, removing these confounds, to address this question and stress test the external validity of model development progress on ImageNet. Additionally, our work focuses on benchmarks as a tool for advancing science; there are many other interpretations of benchmarks that one may be interested in, e.g., as instruments for producing valid scores of a model’s capabilities. Our demonstration of the robustness of rankings under certain benchmark properties, e.g., noise, should not be assumed to extend to other interpretations of benchmarks.
In a nutshell, we give a strong affirmative answer to this question of the external validity. Key model architectures from the ImageNet era rank identically on ImageNot as they do on ImageNet. Moreover, each model’s relative improvement compared with prior models is approximately the same on both datasets. Our empirical findings provide strong evidence for the external validity of model rankings.
@@ 1.1 Our contributions
Figure 1: Final Model Ranking and Relative Improvement. Model rankings and relative improvements hold with models trained from scratch on the respective dataset, and evaluated on a held-out test set. For a set of test accuracies $X$, relative progress is $X_{i}/X_{\text{AlexNet}}$. ImageNet model accuracies range from 0.57 (AlexNet) to 0.85 (EfficientNet V2 L), while much noisier ImageNot model accuracies approximately range from 0.4 (AlexNet) to 0.6 (EfficientNet V2 L).
Figure 2: Examples from ImageNet (left) and ImageNot (right). ImageNot classes (e.g., ‘cleats,’ ‘batter’) are disjoint and distinct from ImageNet classes (e.g., ‘Irish Terrier,’ ‘Blenheim Spaniel’). ImageNot, web-sourced data with automated data selection and labeling, also has greater natural variability and label noise.
@@
Our primary contribution is an in-depth investigation into the external validity, robustness, and reproducibility of model development progress in machine learning. We create a new dataset and perform compute-intensive training and evaluation of numerous architectures on the new dataset.
@@
Creating ImageNot. We created ImageNot from the LAION-5B dataset, a public collection of 5.8 billion image-text pairs crawled from the web and filtered by the CLIP model (schuhmann2022laion; radford2021learning). Although CLIP is trained with knowledge of ImageNet, the sheer size of LAION-5B ensures that images in LAION-5B generally are not similar to ImageNet images. We choose 1000 classes from the WordNet lexical database while avoiding nouns close to those in ImageNet. For each class, we pick 1000 images from LAION-5B by caption-class similarity alone. We use Roberta text embeddings for this purpose. As a text-only model, Roberta was never trained on ImageNet. This mitigates the concern that our selection of images from LAION-5B might be biased in favor of ImageNet-like instances. We give a detailed analysis of the differences between ImageNet and ImageNot. Crucially, our creation of ImageNot avoids confounds in previous work with respect to images and classes, data source, and generative mechanism. We provide details instructions for deriving ImageNot from LAION-5B.
@@
ImageNot preserves model rankings and relative improvements. Our primary empirical finding is that ImageNot has the same image classification model rankings as ImageNet. This finding is important because it provides evidence that the model architecture improvements centered around ImageNet could have occurred similarly for a different dataset than ImageNet. We consider 9 representative model architectures from AlexNet to Vision Transformers (krizhevsky2012imagenet; dosovitskiy2020image). We train from scratch and evaluate each model on ImageNot. We observe that the final ranking is the same as on ImageNet. Moreover, each model’s improvement relative to prior models on ImageNot is roughly the same as on ImageNet. Figure 1 summarizes these results. The remarkable consistency of relative improvements suggests an intriguing counterfactual. Had ImageNot, or other large-scale datasets with different properties, been the benchmark at the time of a new image classification model development, it could have induced the same improvements as ImageNet. Crucially, we train from scratch on ImageNot and evaluate on an ImageNot test set, so we avoid the confound of model training that is pervasive in previous work to truly test external validity.
@@
ImageNot pretraining and transfer learning. ImageNet has proved useful for pretraining and transfer learning (kornblith2019better). We wanted to know if the same is true for ImageNot. Toward this end, we conducted two experiments for each model:
@@
    Pretrain on ImageNot/ImageNet, finetune all layers on CIFAR-10, evaluate on CIFAR-10.
@@
    Pretrain on ImageNot/ImageNet, finetune last layer on CIFAR-10, evaluate on CIFAR-10 – transfer.
Similar to ImageNet, we find that a model’s test accuracy on ImageNot is also a good predictor of fine-tunability and transferability.
To summarize, rankings of image classification models are surprisingly stable across significantly different datasets. Prior work had observed that small variations in a data-generating process preserve model rankings. Our work adds a more extreme contrast. Even significant deliberate deviations from a dataset can maintain essential characteristics of model training and comparison. In particular, our findings suggest that the external validity of machine learning is, in important ways, significantly greater than previously assumed—Code.
@@ 2 Discussion of related work
Early work on dataset bias. The validity of benchmark results for image classification has long been of concern. Indeed, benchmarks often contain biases that do not generalize to data observed in the real world. ponce2006dataset provide evidence of the narrow range of variation within existing benchmarks in the early stages of image classification research. Subsequently, torralba2011unbiased compare multiple well-known image classification datasets, including SUN09 (xiao2010sun), PASCAL Visual Object Classes (VOC) (everingham2010pascal), and Caltech-101 (fei2004learning). They find evidence of selection bias, capture bias, and negative set bias in these datasets. hardtrecht2022patterns Chapter 8 and (liberman2015reproducible) give background and history of datasets and benchmarks.
@@
The ImageNet era. The desire for a more representative benchmark for real-world image classification tasks motivated the creation of ImageNet, a large-scale image database (deng2009imagenet). The extensive scale and careful curation of ImageNet aimed to alleviate some of the limitations identified in earlier datasets. ImageNet was at the center of over a decade of active development in image classification, largely centered around improving accuracy on the ImageNet ILSVRC 2012 competition dataset. The observed progress on image classification can be well summarized by a set of models that span the active development of model architectures around ImageNet and a couple of modern models developed after the challenged ended in 2017: AlexNet (krizhevsky2012imagenet), VGG (simonyan2014very), Inception V3 (szegedy2015going), ResNet (he2016deep), DenseNet (huang2017densely), MobileNet V3 Large (howard2019searching), EfficientNet (tan2019efficientnet), ViT-B-16 (dosovitskiy2020image), and ConvNeXt (liu2022convnet). While other models were also developed during the so-called ImageNet era, we focus on models that were most impactful and represented a notable shift in architecture design—we include a discussion in Appendix A. Specifically, we considered the availability of model implementations across hubs such as TorchHub (pytorch_vision_classification), TensorFlow (abadi2015tensorflow), and Huggingface (wolf2020transformers) to be an indicator of popularity and, therefore, impact.
Investigations about ImageNet. Not least due to its excessive use over the years, researchers registered concerns about the external validity of ImageNet results. pmlr-v119-tsipras20a studied the design choices of ImageNet and their effect on alignment between the ImageNet dataset and the broader underlying object recognition task. They concluded that ImageNet contains some inherent biases that cause misalignment. Despite this misalignment, they argue, improvements in multiclass classification on ImageNet still translate to the underlying object recognition task. However, they raised the issue that these biases make it difficult to identify actual improvements in object recognition. Xiao2020NoiseOS identify one of these biases in that models tend to rely on the background of ImageNet images, which are correlated with the corresponding labels. recht2019imagenet investigated whether ImageNet classifiers generalize to a newly curated ImageNet test set—recall that by ImageNet, we generally refer to the ILSVRC 2012 ImageNet challenge. They found that accuracies decrease strongly from the original the new test set, in the range of 11%-14%. barbu2019objectnet introduce a test set (ObjectNet) with real-world viewpoint, background, and rotation variation to assess model robustness and show that standard ImageNet-trained models suffer a 40–45% drop in top-1 accuracy on ObjectNet. Still, recht2019imagenet found that model rankings did not significantly change between the two ImageNet test sets: if model a performed better than model b on the original ImageNet test set, it generally also performed better on the new ImageNet test set. This speaks to the stability of model rankings under a slight distribution shift (new versus old test set), a form of internal validity. In other words, model rankings replicate under similar testing conditions. This form of internal validity is also supported by studies involving other variants of ImageNet. Hendrycks2019BenchmarkingNN introduced ImageNet-C/P, ImageNet with common corruptions. Hendrycks2020TheMF introduced ImageNet-R, stylized distribution shifts on ImageNet. Li2022AWD introduced ImageNet-W, ImageNet with watermarks, a learning shortcut. shirali2023makes introduced LAIONet, an intended recreation of ImageNet from the LAION-400M dataset, in order to identify key differences between ImageNet and LAIONet.
External validity of ImageNet model rankings. From prior work, we know that model rankings satisfy internal validity, i.e., they replicate in similar conditions, whereas accuracy numbers do not. In this work, we ask a daring question about the external validity of model rankings: Do model rankings replicate in radically different conditions? Rather than studying variants of ImageNet and other smaller and similar datasets to ImageNet, we purposefully create a deliberately different dataset, in a sense, an anti-replication of ImageNet. Surprisingly, we find that models from the ImageNet era achieve the same relative improvements on ImageNot over prior models as they did on ImageNet.
In contrast, we train models from scratch on the new dataset rather than reevaluate ImageNet-trained models on new test sets. One exception, perhaps closest to our study, is the important work of kornblith2019better, who examined how well the performance on ImageNet predicts performance on other vision datasets. They also train from scratch and find that on smaller and older datasets, such as DTD, VOC2007, or Caltech-101, model rankings do not replicate cleanly; however, they also find a strong correlation between ImageNet accuracy and accuracy on these datasets. While kornblith2019better considered standard computer vision datasets at a time, we purposefully create a dataset as different as possible from ImageNet from a data source (LAION) not available then. The key takeaway from  kornblith2019better could have been that ImageNet is special insofar as models developed on ImageNet transfer well. Our main conclusion, in contrast, is that from the perspective of model benchmarking, there may be nothing all that special about ImageNet other than its scale and diversity. In particular, from the benchmarking perspective, our study suggests that we do not even seem to need clean labels (beyer2020we). This surprising takeaway was confirmed in a theoretical study (dorner2024dont).
@@ 3 Constructing ImageNot
Figure 3: Overview of the ImageNot dataset curation pipeline. Starting from the LAION-5B corpus, the process selects all candidate noun synsets, removes classes overlapping with ImageNet or containing too few samples, filters out NSFW or ambiguous data, and ranks remaining classes by semantic (RoBERTa) and visual (CLIP) similarity. The final stage yields 1,000 ImageNot classes matched in scale but disjoint in concept from ImageNet, enabling evaluation of external validity.
Figure 4: ImageNot caption x synset-gloss RoBERTa embeddings similarity for each class. We show some summary statistics of the similarity distribution across the 1000 classes.
@@
We develop ImageNot, a dataset of the same scale as ImageNet’s ILSVRC-2012—1000 classes with 1000 examples per class—but without key design choices that have been thought critical in ImageNet’s effectiveness (ILSVRC15). First, ImageNot has deliberately arbitrary classes constrained only to be strictly different from those in ImageNet. Second, we do not utilize any human annotators. There were no humans in the loop in selecting specific image and label pairs. We create the ImageNot dataset from the LAION-5B dataset, an academic, public massive scale collection of 5 billion image-text pairs crawled from the web and filtered by the CLIP model (schuhmann2022laion; radford2021learning). This contrasts the ILSVRC 2012 challenge, a subset of the ImageNet database which was constructed from Flickr search results (deng2009imagenet).
ImageNot classes correspond to nouns in WordNet synsets found in LAION-5B captions, where WordNet is a large lexical database of English (fellbaum1998WordNet; miller1995WordNet; hardeniya2016natural). The following definitions are important for describing WordNet. A given word has a set of synsets, where a synset is a sense of a word and has a corresponding short definition called a gloss. For example, ‘car’ has 2 synsets with corresponding glosses: (i) “a motor vehicle with four wheels; usually propelled by an internal combustion engine” and (ii) “a wheeled vehicle adapted to the rails of a railroad.” Additionally, each synset consists of a set of synonyms called lemmas. The lemmas of ‘car’-(i) include automobile, motorcar, while the lemmas of ‘car’-(ii) include railcar, railwaycar. A WordNet tree represents a semantic hierarchy of synsets via an ‘IS-A’ structure, i.e., a synset is a ‘type’ of its synset parent, e.g., mammal $\rightarrow$ placental $\rightarrow$ carnivore $\rightarrow$ dog $\rightarrow$ working dog $\rightarrow$ husky.
At a high level, for each class, we select images from LAION-5B whose captions are similar to the synset corresponding to the class. We utilize RoBERTa (liu2019roberta) embeddings, an optimized version of BERT (devlin-etal-2019-bert), to estimate text sequence similarities. Since RoBERTa was never trained on ImageNet, this selection mechanism mitigates the concern that our image selection process may be biased in favor of ImageNet-like images.
We note that creators of LAION-5B’s only included image and caption pairs that are sufficiently similar based on CLIP embeddings (Contrastive Language–Image Pre-training (radford2021learning)), a multimodal deep learning model (schuhmann2022laion).
Constructing ImageNot. In creating ImageNot, we build on the data-generating process developed by shirali2023makes. To choose the 1000 synsets (classes) that are included in ImageNot, available at Code, we take the following steps:
@@
    Filter out synsets considered ‘Not Safe for Work’ (NSFW) by song2021large’s NSFW classifier.
@@
    Filter out synsets with an average top n caption x synset-gloss pair RoBERTa embeddings similarity (normalized inner product) less than 0.82  (liu2019roberta). We found 0.82 to be the highest reasonable choice for sufficient examples in enough classes (Figure 5).
@@
    Finally, select the top 1000 classes ordered by the average similarity between the top 1100 class x image pairs (radford2021learning).
Finally, we obtain ImageNot, a large-scale dataset that contrasts ImageNet. ImageNot has an average caption-label definition similarity of 0.91. Next, we compare ImageNot to ImageNet.
@@
Figure 7: ConvNext distinguishes ImageNot vs. ImageNet images with high accuracy. Error bars are over 3 trials.
@@
In this section, we examine the differences between ImageNot and ImageNet, focusing on their design aspects. The primary distinctions are: (i) the classes in ImageNot are deliberately far and different from those in ImageNet, and (ii) the attribution of a synset to an image in ImageNot is conducted without human intervention. We begin by examining the key differences between the two datasets.
@@
The ‘Not’ in ImageNot. The first significant not is the hierarchy of synsets, based on the WordNet tree. ImageNet is constructed with synsets representing a densely populated subset of the WordNet tree. For instance, ImageNet contains 147 categories of dogs. In contrast, ImageNot is not designed to have densely populated subsets of the WordNet tree. Moreover, ImageNot synsets tend to be higher up in the WordNet tree (Appendix B.1 Figure 11) with ImageNot synsets having a median depth of 8 while ImageNet synsets have a median depth of 10 (more details about the WordNet subtrees can be found in Appendix B Table 3). One consequence of this difference in median depth is that the average diversity of images in an ImageNot synset is higher than that of ImageNet. We illustrate the distribution of intra-class similarity within each dataset in Figure 10, where ImageNet exhibits higher intra-class diversity than ImageNot.